$SPAD/src/input d02raf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
n:=3
mnp:=40
numbeg:=2
nummix:=0
tol:=1.0e-4
init:=0
iy:=3
ijac:=1
lwork:=1925
liwork:=283
np:=17
x:Matrix SF:=
 [[0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,_
   0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,10.0 ,0.0 ,0.0 ,0.0 ,_
   0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,_
   0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ,0.0 ]]
y:Matrix SF:=new(3,40,0.0)
deleps:=0.1
ifail:=11
vef:Vector Expression Float:=
 [Y[2] ,Y[3] ,-Y[1]*Y[3] - 2*EPS*(1-Y[2]*Y[2]) ]
fcn:Asp41(FCN,JACOBF,JACEPS):= retract vef
vei:Vector Expression Integer:=
 [YA[1] ,YA[2] ,YB[2] -1 ]
g:Asp42(G,JACOBG,JACGEP):=retract vei
result:=d02raf(n,mnp,numbeg,nummix,tol,init,iy,ijac,lwork,_
               liwork,np,x,y,deleps,ifail,fcn,g)

———-

References^{[1] nothing}

4